# Axion Icebergs: Clockwork ALPs at hadron colliders

### v1

@@ 2.2 A KSVZ axion
where $Q^{(\alpha)}_{L}$ are the SM quark doublets, $u_{R}^{(\alpha)}$ the up-type singlets and $H$ the SM Higgs doublet. As for the dimensionless couplings $\lambda_{\Psi}$, $y^{(\alpha)}_{\Psi}$ and $y^{{}^{\prime}(\alpha)}_{\Psi}$, the first determines the mass of the $\Psi$ while the other two determine its decay rate. For the sake of simplicity, we assume that $\Psi_{L,R}$ couple to only the third-generation quarks, thereby precluding large flavour-changing neutral currents.
In addition, the CW sector can also couple with the SM Higgs through [^6] For simplicity, we assume uniform value of the coupling $\lambda_{\Phi-H}$ for all $j$. Relaxing this only adds layers of complications without any qualitative change.,
$\mathcal{L}_{\Phi-H}=-\lambda_{\Phi H}\sum_{j=0}^{N}\Phi_{j}^{\dagger}\Phi_{j}H^{\dagger}H$ (9)
@@
[…]
A rephasing of the new fermions with respect to the Goldstone field $\pi_{N}$ leads to the following interaction terms with the SM gauge bosons via the chiral anomaly (see appendix A for a discussion)
@@
2.2.1 The Peccei-Quinn Mechanism
In order for the light pseudoscalar to be the quintessential QCD axion, it must pose a solution to the strong CP problem. Although ref.Farina:2016tgd already has a brief discussion in this context, we outline here, for completeness, how the clockwork Lagrangian can be consistent with the usual Peccei-Quinn mechanism. The foremost requirement for the mechanism is the presence of a global symmetry (atleast an accidental one) in the theory. In our case this symmetry would correspond to the residual clockwork symmetry described in the preceding section (see eq.(3)). Now, in eq.(11), we can perform a shift in the $\pi_{N}$ field so as to exactly cancel the CP violating topological term $-\bar{\theta}(\alpha_{S}/8\pi)\tilde{G}^{\mu\nu}G_{\mu\nu}$. This, however, is a true cancellation, and a potential solution to the strong CP problem, only when $\pi_{N}$ does not acquire a nonzero vacuum expectation value (VEV) after employing the constant shift. It is important to note that the explicit breaking of the $[U(1)]^{N+1}$ symmetry in the clockwork sector, courtesy the nearest-neighbour interactions, does not generate a nonzero VEV for the pseudoscalar fields thanks to the residual shift symmetry $U(1)_{CW}$ Banerjee:2018grm . The only other source from where $\pi_{N}$ gets a potential is due to nonperturbative effects below the QCD confinement scale. This potential would have the approximate form,
$V(\pi_{N})\sim\Lambda^{4}_{QCD}\left[1-\cos{\left(\frac{\pi_{N}}{f_{\rm eff}}\right)}\right],$ (13)
which shows that $\pi_{N}$ does not, indeed, acquire a nonzero VEV [^7] Naively, the potential would have a degenerate set of minima at $\langle\pi_{N}\rangle=2n\pi f_{\rm eff}$ for integral values of $n$. However, such nonzero VEVs are unphysical as they can always be shifted to zero by the residual symmetry transformation $\pi_{N}\to\pi_{N}+2n\pi f_{\rm eff}$ as a result of the $2\pi$ periodicity of the potential.. Going to the mass basis of the CW axions, the lightest axion, i.e. the KSVZ QCD axion, would then obtain a mass given by Alonso-Alvarez:2018irt ; GrillidiCortona:2015jxo ,
@@ 2.2.1 The Peccei-Quinn Mechanism
It is perhaps useful to remark here that the potential in eq.(13) is a special case which defines a unique vacuum with a domain wall number $N_{DW}=1$. This is a direct consequence of the fact that we have invoked only a single generation of the extra $SU(2)$ singlet coloured fermion $\Psi$. Similar to the $N_{DW}>1$ scenarios, cosmological domain walls would also form in this case at temperatures near the QCD phase transition scale, albeit with a simple structure where each wall is bounded by an individual cosmic string. Such a simple string-wall network, however, is highly unstable and would be prone to a rapid collapse Barr:1986hs ; Kim:1986ax ; Chang:1998tb . Thus, apart from the obvious simplicity, the choice of a single species of the singlet heavy quark possesses the important feature of alleviating the domain wall problem Sikivie:1982qv typical to axion models.
Figure 1: Limits on the effective decay constant $f_{0}$ of the light QCD axion vs its mass. Plot adapted from ref.AxionLimits .
2.3 Axion physics above the QCD confinement scale
@@
$\mathcal{L}_{\pi vv}=-g_{\pi gg}\,\pi_{N}G^{A\mu\nu}\tilde{G}^{A}_{\mu\nu}-g_{\pi\gamma\gamma}\,\pi_{N}F^{\mu\nu}\tilde{F}_{\mu\nu}-g_{\pi\gamma Z}\,\pi_{N}F^{\mu\nu}\tilde{Z}_{\mu\nu}-g_{\pi ZZ}\,\pi_{N}Z^{\mu\nu}\tilde{Z}_{\mu\nu}.$ (15)
The coefficients in this case are given by,
$\begin{array}[]{rclcrcl}g_{\pi gg}&=&\displaystyle\frac{\alpha_{s}}{8\pi f_{\rm eff}}\,,&&g_{\pi\gamma\gamma}&=&\displaystyle\frac{2N_{c}\alpha_{EM}Y^{2}_{\Psi}}{8\pi f_{\rm eff}}\,,\\[8.61108pt] g_{\pi\gamma Z}&=&\displaystyle\frac{-4N_{c}s^{2}_{w}\alpha_{EM}Y^{2}_{\Psi}}{8\pi f_{\rm eff}s_{w}c_{w}}\,,&&g_{\pi ZZ}&=&\displaystyle\frac{2N_{c}s^{4}_{w}\alpha_{EM}Y^{2}_{\Psi}}{8\pi f_{\rm eff}s_{w}^{2}c_{w}^{2}}.\end{array}$ (16)
where $s_{w}\equiv\sin\theta_{w}$ and $c_{w}\equiv\cos\theta_{w}$ with $\theta_{w}$ being the Weinberg angle. The effective couplings of the physical axions are then readily obtained using eq.(6),
$\begin{array}[]{rcl c rcl}|g^{(n)}_{agg}|&=&\displaystyle\frac{\alpha_{s}}{8\pi f_{n}}\,,&&|g^{(n)}_{\pi\gamma\gamma}|&=&\displaystyle\frac{2N_{c}\alpha_{EM}Y^{2}_{\Psi}}{8\pi f_{n}}\,,\\[8.61108pt] |g^{(n)}_{a\gamma Z}|&=&\displaystyle\frac{4N_{c}s^{2}_{w}\alpha_{EM}Y^{2}_{\Psi}}{8\pi f_{n}s_{w}c_{w}}\,,&&|g^{(n)}_{\pi ZZ}|&=&\displaystyle\frac{2N_{c}s^{4}_{w}\alpha_{EM}Y^{2}_{\Psi}}{8\pi f_{n}s_{w}^{2}c_{w}^{2}}\end{array}$ (17)
@@
where $f_{n}\equiv f_{\rm eff}/|C_{nN}|$ denotes the effective decay constant of the $n$-th physical axion. It is then readily apparent that the decay constants of the heavy pseudoscalars are hierarchically smaller than that of the light axion. For a SSB scale $f\sim\mathcal{O}(\mbox{TeV})$, this engenders significantly enhanced couplings of the heavy axions with gluons and the electroweak bosons as compared with those associated with typical light ALP candidates. This enhancement, in principle, can facilitate resonant production of the heavy CW axions at hadron colliders, which is going to be the main subject of discussion in the ensuing sections.
3 Clockwork axions at colliders
@@
As is already described above, the (light) axion, when compatible with the constraints from astrophysical and low-energy experiments Semertzidis:2021rxs , is nearly invisible at high energy collider experiments. What seems promising, instead, is the prospect of probing the heavy pseudoscalars (hereafter, also referred to as axion-like-particles or ALPs) at the LHC and its forthcoming high-luminosity upgrade (HL-LHC). In this case, the production would be dominated by the gluon fusion channel with the hadronic cross-section given by
$\hat{\sigma}(pp\to a_{n})=K^{0}_{\sigma}\frac{d\mathcal{L}_{gg}}{d\hat{s}}\Bigg{|}_{\hat{s}=m_{n}^{2}}\frac{\pi^{2}}{8m_{n}}\Gamma^{(n)}_{gg}\delta(\hat{s}-m_{n}^{2}),$ (18)
@@
where $d\mathcal{L}_{gg}/d\hat{s}$ is the gluon-gluon luminosity and the $K$-factor $K^{0}_{\sigma}$ encapsulates the higher order QCD corrections. Using the MSTW2008nnlo68 parton densities, ref.Mariotti:2017vtv estimates $K^{0}_{\sigma}\approx 3.7$. (including the full NNLO and approximate N3LO corrections Ball:2013bra ; Bonvini:2014jma ; Bonvini:2016frm ; Ahmed:2016otz ) over the pseudoscalar mass range $40-125$ GeV . In the absence of a full computation of the QCD corrections, we make the reasonable extrapolation that $K^{0}_{\sigma}\sim 3.7$ for the mass range $10-125$ GeV. For ALP masses beyond $150$ GeV we assume a more conservative $K^{0}_{\sigma}=2.5$ Ahmed:2016otz .
@@
$\begin{split}\Gamma(a_{n}\to gg)\equiv\Gamma^{(n)}_{gg}&=K_{gg}C_{n,N}^{2}|g(x_{n})|^{2}\frac{\alpha_{S}^{2}}{32\pi^{3}}\frac{m_{n}^{3}}{f_{\mbox{eff}}^{2}}\\ \Gamma(a_{n}\to\gamma\gamma)\equiv\Gamma^{(n)}_{\gamma\gamma}&=9Y_{\Psi}^{4}C_{n,N}^{2}|g(x_{n})|^{2}\frac{\alpha_{EM}^{2}}{64\pi^{3}}\frac{m_{n}^{3}}{f_{\mbox{eff}}^{2}}\\ \Gamma(a_{n}\to Z\gamma)\equiv\Gamma^{(n)}_{Z\gamma}&=9Y_{\Psi}^{4}C_{n,N}^{2}|g(x_{n})|^{2}\frac{\alpha_{EM}^{2}\tan^{2}{\theta_{w}}}{8\pi^{3}}\frac{(m_{n}^{2}-m_{Z}^{2})^{3}}{m_{n}^{3}f_{\mbox{eff}}^{2}}\\ \Gamma(a_{n}\to ZZ)\equiv\Gamma^{(n)}_{ZZ}&=9Y_{\Psi}^{4}C_{n,N}^{2}|g(x_{n})|^{2}\frac{\alpha_{EM}^{2}\tan^{4}{\theta_{w}}}{64\pi^{3}}\,\frac{(m_{n}^{2}-4m_{Z}^{2})^{3/2}}{f_{\mbox{eff}}^{2}}\end{split},$ (19)
where $x_{n}\equiv 4m_{\Psi}^{2}/m_{n}^{2}$ and the loop-integral $g(x)$ is defined as,
@@
In the expression for $\Gamma_{gg}$, the QCD correction amounts to $K_{gg}=2.1$ Djouadi:2005gi . Clearly, the ALPs decay dominantly to two gluons followed by decays to photons (and to $Z\gamma$ and $ZZ$ depending on kinematic feasibility). It is easy to infer in the narrow width approximation that the dijet cross-section would exceed the diphoton one by a factor of $[2/(9Y_{\Psi}^{4})]K_{gg}(\alpha_{S}^{2}/\alpha_{EM}^{2})$. However, as far as detection at hadron colliders is concerned, the diphoton final state, understandably, offers better sensitivity as compared to the dijet channel. We will, therefore, consider only the processes $pp\to a_{n}\to\gamma\gamma$ for the case in hand.
[…]
@@
For each of the benchmark points, we choose $\xi=3$ and $q=2$ (as the middle point of the theoetically allowed range) and also assume $Y_{\Psi}=2/3$ so that the $\Psi$ may mix with the top quark. The other particulars are as follows—
Benchmark I: We assume the parameter values $N=28$, $m=10$ GeV and $f=1000$ GeV in the CW sector. The resulting light axion mass in this case is $m_{a_{0}}\sim 6.4\times 10^{-5}$ eV and the ALP masses span the range $\sim 10-30$ GeV. This particular choice of the CW parameters, especially the sizable number of particles $N$, is also motivated by the fact that this allows for the light axion to be a DM candidate Alonso-Alvarez:2023wig . With these choices in place, we can determine the diphoton cross-section in the narrow-width approximation (NWA). Fig.2(a) shows the cross-sections for each of the ALPs (integrated over the full final state phase-space) computed for a center of mass energy $\sqrt{s}=13$ TeV using the NNPDF2.3LO PDF set NNPDF:2014otw . Note that the cross-section peaks for an intermediate mode of the spectrum, the exact identity of which depends on the periodic variation of the couplings due to the clockwork mixing as well as the energy dependence of the parton-parton luminosity. Fig.2(b), on the other hand, shows $\Delta m_{n}$ (the mass-splitting between consecutive CW modes) along the full spectrum, with the average mass-splitting being given by $\Delta m\sim 2m/N$. It is interesting to note that, with the consecutive differences $\Delta m_{n}\lesssim 1$ GeV over the entire extent of the ALP spectrum, the splittings are comparable with the prevailing detector resolution. Consequently, the individual resonances may not be entirely resolved and the events corresponding to the full spectrum (subject to the assumed bin size) may even appear as a single broad resonance [^8] This was also hinted at in ref.Higaki:2015jag . within an envelope of mass width $\sim\Delta m$ in the diphoton invariant mass distribution.
@@
Figure 2: (a) Masses and couplings for individual ALPs and corresponding diphoton cross-sections for benchmark I. The vertical axis on the right shows the extent of the field $\pi_{N}$ contained in the mass eigenstates $a_{n}$. (b) Mass-splittings between consecutive ALPs for benchmark I.
@@
Benchmark II: (a) The parameter configuration in the CW sector for this case is — $m=35$ GeV, $N=28$ and $f=1000$ GeV. The light axion mass remains the same as in benchmark I while the ALP masses now range from $35$ to $105$ GeV. The corresponding diphoton cross-sections are shown in Fig.3(a). Due to an increase in the value of the parameter $m$ (by a factor of 3.5) the mass-splittings in this case, readily inferred from eq.(5), are nearly triple in value compared to that in the preceding case. Therefore, in stark contrast with benchmark I, one expects to discern at least a few individual peaks in the invariant mass distribution of the signal events. We will see that this is indeed the case.
Figure 3: Masses and couplings for individual ALPs and corresponding diphoton cross-sections. (Left) benchmark II(a) and (right) benchmark II(b).

### v2

@@ II.2 A KSVZ axion
where $Q^{(\alpha)}_{L}$ are the SM quark doublets, $u_{R}^{(\alpha)}$ the up-type singlets and $H$ the SM Higgs doublet. As for the dimensionless couplings $\lambda_{\Psi}$, $y^{(\alpha)}_{\Psi}$ and $y^{{}^{\prime}(\alpha)}_{\Psi}$, the first determines the mass of the $\Psi$ while the other two determine its decay rate. For the sake of simplicity, we assume that $\Psi_{L,R}$ couple to only the third-generation quarks, thereby precluding sizable flavour-changing neutral currents involving the first two generations.
In addition, the CW sector can also couple with the SM Higgs through [^9] For simplicity, we assume uniform value of the coupling $\lambda_{\Phi-H}$ for all $j$. Relaxing this only adds layers of complications without any qualitative change.
$\mathcal{L}_{\Phi-H}=-\lambda_{\Phi H}\sum_{j=0}^{N}\Phi_{j}^{\dagger}\Phi_{j}H^{\dagger}H\,.$ (9)
@@
$\mathcal{L}_{\Psi}\supset-\frac{1}{\sqrt{2}}\lambda_{\Psi}(\phi_{N}+f)e^{i\xi\pi_{N}/f}\bar{\Psi}_{L}\Psi_{R}-y_{\Psi}\bar{Q}^{(3)}_{L}\tilde{H}\Psi_{R}-\frac{1}{\sqrt{2}}y^{\prime}_{\Psi}(\phi_{N}+f)e^{i\xi\pi_{N}/f}\bar{\Psi}_{L}u^{(3)}_{R}+\mbox{h.c.}\,.$ (10)
A rephasing of the new fermions with respect to the Goldstone field $\pi_{N}$ leads to the following interaction terms with the SM gauge bosons via the chiral anomaly Alonso-Álvarez et al. (2019) (see appendix A for a discussion)
@@
II.2.1 The Peccei-Quinn Mechanism
In order for the light pseudoscalar to be the quintessential QCD axion, it must offer a solution to the strong CP problem. Although ref.Farina et al. (2017) already has a brief discussion in this context, we outline here, for completeness, how the clockwork Lagrangian can be consistent with the usual Peccei-Quinn mechanism. The foremost requirement for the mechanism is the presence of a global symmetry (at least an accidental one) in the theory. In our case, this symmetry would correspond to the residual clockwork symmetry described in the preceding section (see eq.(3)). Now, in eq.(11), we can perform a shift in the $\pi_{N}$ field so as to exactly cancel the CP violating topological term $\bar{\theta}(\alpha_{S}/8\pi)\tilde{G}^{\mu\nu}G_{\mu\nu}$. This, however, is a true cancellation, and a potential solution to the strong CP problem, only when $\pi_{N}$ does not acquire a nonzero vacuum expectation value (VEV) after employing the constant shift. It is important to note that the explicit breaking of the $U(1)]^{N+1}$ symmetry in the clockwork sector, courtesy the nearest-neighbour interactions, does not generate a nonzero VEV for the pseudoscalar fields thanks to the residual shift symmetry $U(1)_{CW}$ Banerjee et al. (2018). The only other source from where $\pi_{N}$ may acquire a nontrivial potential are nonperturbative effects below the QCD confinement scale, leading to an effective potential of the form Di Vecchia and Veneziano (1980); Grilli di Cortona et al. (2016)
$V(\pi_{N})=-m_{\pi}^{2}f_{\pi}^{2}\sqrt{1-\frac{4m_{u}m_{d}}{(m_{u}+m_{d})^{2}}\sin^{2}{\left(\frac{\pi_{N}}{2f_{\rm eff}}\right)}}$ (13)
where $f_{\pi}$ and $m_{\pi}$ are the pion decay constant and mass, respectively, and $m_{u,d}$ are the masses of the light SM quarks $u$ and $d$. This shows that $\pi_{N}$ does not, indeed, acquire a nonzero VEV [^10] Naively, the potential would have a degenerate set of minima at $\langle\pi_{N}\rangle=2n\pi f_{\rm eff}$ for integral values of $n$. However, such nonzero VEVs are unphysical as they can always be shifted to zero by the residual symmetry transformation $\pi_{N}\to\pi_{N}+2n\pi f_{\rm eff}$ as a result of the $2\pi$ periodicity of the potential.. Expanding the potential in powers of $\pi_{N}$, the quadratic term engenders a mass correction to $\pi_{N}$
$\Delta m_{\pi_{N}}^{2}=\frac{f_{\pi}^{2}m_{\pi}^{2}}{f_{\rm eff}^{2}}\frac{m_{u}m_{d}}{(m_{u}+m_{d})^{2}}\,.$ (14)
Going to the physical basis of the CW axions, this generates a small mass for the lightest state, i.e., the KSVZ QCD axion, given by
@@ II.2.1 The Peccei-Quinn Mechanism
Figure 1: Limits on the effective decay constant $f_{0}$ of the light QCD axion vs its mass. Plot adapted from ref.O’Hare (2020).
It is perhaps prudent to remark at this juncture about the possible nature and implications of the topological defects that may arise in the multi-axion configuration of the clockwork sector. The spontaneous breaking of the $U(1)^{N+1}$ symmetry at a temperature $T\sim f\gg\Lambda_{QCD}$ (post inflation and reheating) is expected to lead to the formation of $N+1$ cosmic strings bound to one another through domain walls Long (2018); Higaki et al. (2016b); Lee et al. (2024). Such a string-wall network would, in principle, tend to radiate light axions as well as ALPs (as also gravitational waves Higaki et al. (2016b); Lee et al. (2024)) during the course of its cosmological evolution towards the epoch of the QCD phase transition [^11] The ALPs, owing to their relatively small decay constant, decay promptly once they are produced.. In a canonical framework, these emissions are estimated to scale as $f^{2}$ and, therefore, constitute a large fraction of the axion abundance in the usual QCD axion cosmology where $f\sim f_{0}\gtrsim 10^{9}\,\mbox{GeV}$. In the present context, though, the clockwork mechanism generates an exponential separation between the SSB scale $f$ and the effective decay constant of the light axion $f_{0}$, such contributions to the axion abundance is exceedingly suppressed for $f\sim\mathcal{O}(\mbox{TeV})$. At the epoch of the QCD phase transition, the induced potential of eq.(13) defines a unique vacuum for the system associated with a domain wall number $N_{DW}=1$ — a direct consequence of invoking only a single generation of the extra $SU(2)$ singlet coloured fermion $\Psi$. Similar to the case of a single axion Barr et al. (1987); Kim (1987); Chang et al. (1999), such a potential destabilizes the string-wall network and leads to its eventual collapse [^12] Thus, apart from the obvious simplicity, the choice of a single species of the singlet heavy quark possesses the important feature of alleviating the domain wall problem Sikivie (1982) typical to axion models., with its energy being released primarily in the form of axions and ALPs. This contribution to the axion abundance, too, can be estimated to be insignificant, with the axion emission being suppressed by the factor $\sim(m_{a_{0}}/m)^{3}$ Long (2018).
While a certain abundance of the axion as dark matter is expected to be generated via radiation and collapse of the topological defects in the clockwork system, a robust estimate of the same can only be obtained from a dedicated numerical simulation which is extremely complicated to perform for a multi-axion setup, especially for $N\gg\mathcal{O}(1)$. In the absence of such analyses, we admit to a reasonable assertion, based on the aforementioned arguments, that for the SSB scale of interest $f\sim\mathcal{O}(\mbox{TeV})\ll f_{0}$ the dominant contribution to the axion DM abundance emanates from the misalignment mechanism Fox et al. (2004), namely
$\Omega h^{2}\simeq 0.2\left(\frac{f_{0}}{10^{11}{\rm GeV}}\right)^{7/6}\left(\frac{\theta_{0}^{2}}{\pi^{2}/3}\right)$ (16)
where $\theta_{0}$ is the average initial misalignment angle of the QCD axion. This implies that for the natural choice of $\theta_{0}\sim\mathcal{O}(1)$ in eq.(16), the constraint from overclosure demands $f_{0}\lesssim 10^{11}\,\mbox{GeV}$. The experimental and observational limits demand, on the other hand, $f_{0}\gtrsim 10^{9}\,\mbox{GeV}$. The allowed window is then wide enough and comparable with the one usually encountered for the traditional QCD axion. Indeed, this is why it is also the target region for several cavity haloscope experiments such as ADMX, CAPP, HAYSTAC, QUAX, etc. (see e.g. figure 90.5 of the PDG review on axions Navas et al. (2024)).
In the spirit of these arguments, we would be assuming, in the discussions to follow, benchmarks that are consistent with the above mentioned bounds on $f_{0}$.
II.3 Axion physics above the QCD confinement scale
@@
$\begin{split}\mathcal{L}_{\pi vv}=&-g_{\pi gg}\,\pi_{N}G^{A\mu\nu}\tilde{G}^{A}_{\mu\nu}-g_{\pi\gamma\gamma}\,\pi_{N}F^{\mu\nu}\tilde{F}_{\mu\nu}\\ &-g_{\pi\gamma Z}\,\pi_{N}F^{\mu\nu}\tilde{Z}_{\mu\nu}-g_{\pi ZZ}\,\pi_{N}Z^{\mu\nu}\tilde{Z}_{\mu\nu}.\end{split}$ (17)
The coefficients in this case are given by
[…]
where $s_{w}\equiv\sin\theta_{w}$ and $c_{w}\equiv\cos\theta_{w}$ with $\theta_{w}$ being the Weinberg angle. For the mass-eigenstates, the couplings are simply scaled by the mixing parameters, viz.
$g^{(n)}_{agg}=g_{\pi gg}(f_{\rm eff}\to f_{n})$ (18)
@@
where $f_{n}\equiv f_{\rm eff}/C_{nN}$ denotes the effective decay constant of the $n$-th physical axion. Using eq.(6), it is then readily apparent that the decay constants of the heavy pseudoscalars are hierarchically smaller than that of the light axion. For a SSB scale $f\sim\mathcal{O}(\mbox{TeV})$, this engenders significantly enhanced couplings of the heavy axions with gluons and the electroweak bosons as compared with those associated with typical light ALP candidates. This enhancement, in principle, can facilitate resonant production of the heavy CW axions at hadron colliders, which is going to be the main subject of discussion in the ensuing sections.
III Clockwork axions at colliders
@@
As is already described above, the (light) axion, when compatible with the constraints from astrophysical and low-energy experiments Semertzidis and Youn (2022), is nearly invisible at high energy collider experiments. What seems promising, instead, is the prospect of probing the heavy pseudoscalars at the LHC and its forthcoming high-luminosity upgrade (HL-LHC). In this case, the production would be dominated by the gluon fusion channel with the hadronic cross-section given by
$\sigma(pp\to a_{n})=K^{0}_{\sigma}\frac{d\mathcal{L}_{gg}}{d\hat{s}}\Bigg{|}_{\hat{s}=m_{n}^{2}}\frac{\pi^{2}}{8m_{n}}\hat{\Gamma}^{(n)}_{gg}$, (19)
@@
where $d\mathcal{L}_{gg}/d\hat{s}$ is the gluon-gluon luminosity and the $K$-factor $K^{0}_{\sigma}$ encapsulates the higher order QCD corrections. Using the MSTW2008nnlo68 parton densities, ref.Mariotti et al. (2018) estimates $K^{0}_{\sigma}\approx 3.7$. (including the full NNLO and approximate N3LO corrections Ball et al. (2013); Bonvini et al. (2014, 2016); Ahmed et al. (2016)) over the pseudoscalar mass range 4012540-125 GeV . In the absence of a full computation of the QCD corrections, we make the reasonable extrapolation that $K^{0}_{\sigma}\sim 3.7$ for the mass range 1012510-125 GeV. For ALP masses beyond $150$ GeV we assume a more conservative $K^{0}_{\sigma}=2.5$ Ahmed et al. (2016).
@@
$\begin{split}\Gamma(a_{n}\to gg)&\equiv K_{gg}\hat{\Gamma}^{(n)}_{gg}=K_{gg}|C_{nN}g(x_{n})|^{2}\frac{\alpha_{S}^{2}}{32\pi^{3}}\frac{m_{n}^{3}}{f_{\mbox{eff}}^{2}}\\ \Gamma(a_{n}\to\gamma\gamma)&\equiv\Gamma^{(n)}_{\gamma\gamma}=9Y_{\Psi}^{4}|C_{nN}g(x_{n})|^{2}\frac{\alpha_{EM}^{2}}{64\pi^{3}}\frac{m_{n}^{3}}{f_{\mbox{eff}}^{2}}\\ \Gamma(a_{n}\to Z\gamma)&\equiv\Gamma^{(n)}_{Z\gamma}=9Y_{\Psi}^{4}|C_{nN}g(x_{n})|^{2}\frac{\alpha_{EM}^{2}\tan^{2}{\theta_{w}}}{8\pi^{3}}\frac{(m_{n}^{2}-m_{Z}^{2})^{3}}{m_{n}^{3}f_{\mbox{eff}}^{2}}\\ \Gamma(a_{n}\to ZZ)&\equiv\Gamma^{(n)}_{ZZ}=9Y_{\Psi}^{4}|C_{nN}g(x_{n})|^{2}\frac{\alpha_{EM}^{2}\tan^{4}{\theta_{w}}}{64\pi^{3}}\,\frac{(m_{n}^{2}-4m_{Z}^{2})^{3/2}}{f_{\mbox{eff}}^{2}}\end{split},$ (20)
where $x_{n}\equiv 4m_{\Psi}^{2}/m_{n}^{2}$ and the loop-integral $g(x)$ is defined as
@@
In the expression for $\Gamma(a_{n}\to gg)$, the QCD correction amounts to $K_{gg}=2.1$ Djouadi (2008). Clearly, the ALPs decay dominantly to two gluons followed by decays to photons (and to $Z\gamma$ and $ZZ$ depending on kinematic feasibility). It is easy to infer—within the narrow width approximation— that the dijet cross-section would exceed the diphoton one by a factor of  [^13] Note that $K_{\gamma\gamma}\approx 1$. $2/(9Y_{\Psi}^{4})]K_{gg}(\alpha_{S}^{2}/\alpha_{EM}^{2})$. However, as far as detection at hadron colliders is concerned, the diphoton final state, understandably, offers better sensitivity as compared to the dijet channel. We will, therefore, consider only the processes $pp\to a_{n}\to\gamma\gamma$ for the case in hand.
The diphoton final state has been extensively studied in the literature in the context of spin-0 and spin-2 resonances, with dedicated searches by the ATLAS and CMS collaborations spanning almost the entire range of masses accessible with current sensitivities. With measurements thus far reporting, at best, only marginal excess of events over the SM background, the most recent analyses from ATLAS Aad et al. (2023a, 2021, 2024b) and CMS CMS (2023, 2024) place the strongest limits to date on the $pp\to\gamma\gamma$ fiducial cross-section for spin-0 resonances over a wide mass bracket ranging from nearly $10\,\mbox{GeV}$ to a few TeVs. Being optimistic that some new physics indeed exists beyond the SM, this broadly implies two possibilities —first, that the energy scale of new physics lies beyond the reach of the LHC and, second, that new dynamics exist within the LHC’s energy reach, albeit with feeble interaction strengths with the SM so as to be inaccessible with the current luminosity reach. In the following, we choose the latter viewpoint and explore viable scenarios where the diphoton cross-sections for the CW pseudoscalar spectrum could be very close to the current exclusion limits at the LHC obtained from the Run-2 data, such that a sizable enhancement in the signal significance can be achieved by the end of the ongoing Run-3 phase. This is interesting because even if a $5\sigma$ discovery looks improbable at the LHC, such enhancements could hint at a potential discovery at the HL-LHC taking into account the projected integrated luminosity of $\sim 3000\,\mbox{fb}^{-1}$, i.e. nearly 20 times the luminosity achieved by the end of LHC Run-2 ($\sim 138\,\mbox{fb}^{-1}$). We consider for our study three benchmark points categorized by the ALP masses [^14] While we do not consider ALP masses below $10$ GeV, collider searches for lighter axions in general have been studied in refs.Cid Vidal et al. (2019); Acanfora et al. (2024). in the range 103010-30 GeV, 3510535-105 GeV (with two sub-categories in this case) and 150450150-450 GeV, respectively, and have as guiding references the ATLAS analyses carried out in refs. Aad et al. (2023a, 2021, 2024b).
@@
For each of the benchmark points, we choose $\xi=3$ and $q=2$ (as the middle point of the theoretically allowed range) and also assume $Y_{\Psi}=2/3$ so that the $\Psi$ may mix with the top quark. The other particulars are as follows—
@@
Figure 2: (top) Masses and couplings for individual ALPs and corresponding diphoton cross-sections for benchmark I. The vertical axis on the right shows the extent of the field $\pi_{N}$ contained in the mass eigenstates $a_{n}$. (bottom) Mass-splittings between consecutive ALPs for benchmark I.
Benchmark I: We assume the parameter values $N=28$, $m=10$ GeV and $f=1000$ GeV in the CW sector. The resulting light axion in this case has a decay constant $f_{0}\approx 9\times 10^{10}$ GeV and a mass $m_{a_{0}}\approx 6.4\times 10^{-5}$ eV. The corresponding ALP masses span the range $\sim 10-30$ GeV. This particular choice of the CW parameters, especially the sizable number of particles $N$, is also motivated by the fact that it is compatible with the light axion being a DM candidate [^15] The notion of the KSVZ axion as dark matter has also been explored in ref.Alonso-Álvarez et al. (2023), albeit in the traditional one scalar setup. (with an acceptable relic density and consistency with cosmological observations)—as discussed in section II.2.1—for $\mathcal{O}(1)$ values of the misalignment angle [^16] See figures 4 and 5 in ref.Long (2018) for a summary of the various cosmological constraints on the CW parameter space.. With these choices in place, we can determine the diphoton cross-section in the narrow-width approximation (NWA). Fig.2 (top) shows the cross-sections for each of the ALPs (integrated over the full final state phase-space) computed for a center of mass energy $\sqrt{s}=13$ TeV using the NNPDF2.3LO PDF set Ball et al. (2015). Note that the cross-section peaks for an intermediate mode of the spectrum, the exact identity of which depends on the periodic variation of the couplings due to the clockwork mixing as well as the energy dependence of the parton-parton luminosity. Fig.2 (bottom), on the other hand, shows $\Delta m_{n}$ (the mass-splitting between consecutive CW modes) along the full spectrum, with the average mass-splitting being given by $\Delta m\sim 2m/N$. It is interesting to note that, with the consecutive differences $\Delta m_{n}\lesssim 1$ GeV over the entire extent of the ALP spectrum, the splittings are comparable with the prevailing detector resolution. Consequently, the individual resonances may not be entirely resolved and the events corresponding to the full spectrum (subject to the assumed bin size) may even appear as a single broad resonance [^17] This was also hinted at in ref.Higaki et al. (2016a). within an envelope of mass width  $\sim\,2m$ in the diphoton invariant mass distribution.
@@
Benchmark II(a): The parameter configuration in the CW sector for this case is — $m=35$ GeV, $N=28$ and $f=1000$ GeV. The light axion has the same decay constant and mass as in benchmark I and, thus, is still a potential DM candidate, while the ALP masses now range from $35$ to $105$ GeV. The corresponding diphoton cross-sections are shown in Fig.3. Due to an increase in the value of the parameter $m$ (by a factor of 3.5) the mass-splittings in this case, readily inferred from eq.(5), are nearly triple in value compared to that in the preceding case. Therefore, in stark contrast with benchmark I, one expects to discern at least a few individual peaks in the invariant mass distribution of the signal events. We will see that this is indeed the case.